### v2

@@ Abstract
In this work we propose and analyse a structure-preserving approximation of the non-isothermal Cahn-Hilliard-Navier-Stokes system using conforming finite elements in space and implicit time discretisation with convex-concave splitting. The system is first reformulated into a variational form which reveal the structure of the equations, which is then used in the subsequent approximation.
@@ 1 Introduction
The non-isothermal Cahn-Hilliard-Navier-Stokes (CHNST) system has increasingly gained attention for the investigation of different phenomena ranging from two-phase flows to fluid-phase-coupled interactions both of which have relevant importance in scientific and industrial applications, such as additive manufacturing and inkjet printing van2017binary ; Yang2020 ; dadvand2021advected . For instance, the modeling and simulation of powder bed fusion additive manufacturing (PBF-AM) process employs the non-isothermal CHNST to depict coupled processes in PBF-AM such as fluid-phase interaction, melt flow dynamics and heat transfer Yang2020 . For the CHNST system we consider the following system of partial differential equations
@@
The above system is complemented by periodic boundary conditions and initial conditions. In this context $\phi$ denotes a conserved phase-field, $\mathbf{u}$ is the flow velocity, $\theta$ is the inverse temperature and $e\equiv e(\phi,\theta)$ is the internal energy. To close the system we consider the Helmholtz free energy density and the Korteweg stress given by
@@
$\tilde{\Psi}(\phi,\theta)=\Psi(\phi,\theta)+\tfrac{\gamma}{2}|\nabla\phi|^{2},\qquad\boldsymbol{\sigma}:=\tfrac{\gamma}{\theta}\nabla\phi\otimes\nabla\phi.$
@@
From this, one can compute the internal energy and entropy according to Alt1990 ; Alt1992 , i.e. $e=\partial_{\theta}\tilde{\Psi}$ and $\hat{s}(e(\theta,\phi),\phi)=s(\theta,\phi)=\theta e-\tilde{\Psi}$.
@@
The above system is already formulated in the inverse temperature $\theta$ which is more convenient for a finite element discretisation. Transformation from the original temperature to the inverse temperature can be found by Pawlow and Alt, see Alt1992 , and also brunk2023variational for a more complicated non-isothermal Cahn-Hilliard-Allen-Cahn system. The general derivation of non-isothermal phase-field models from thermodynamic sound principles is more delicate and we refer to Charach1998 ; Fabrizio2006 ; Pawlow2016 for several different modelling approaches.
@@
For such kind of systems there are works Guo_2015 ; Sun2020 , where they consider finite differences in space together with an Energy Quadratisation (EQ) ansatz. This renders the driving functional quadratic, which facilities standard time discretisation methods. However, the price of such an ansatz is the relaxation of entropy, which in some cases results in a drift-off from the original entropy. The isothermal version of the above system, i.e. the Cahn-Hilliard-Navier-Stokes system, is typically called model H and was introduced by Hohenberg and Halperin Hohenberg . This model is employed in the context of two-phase flow especially in phase separation. In this case there are many well-established techniques to solve the system with finite elements Diegel2017 ; Feng06 ; Han2015 and finite differences Gong2018 ; ChenZhao2020 ; LiShen2022 in space.
@@
In this work we will provide a corresponding fully discrete method for the Cahn-Hilliard-Navier-Stokes system using standard finite elements in space and suitable explicit and implicit time-stepping. The discretisation can be seen as an natural extension of the well-known convex-concave splitting for the isothermal Cahn-Hilliard-Navier-Stokes system to the non-isothermal case.
@@
The work is structured as follows. In Section 2 we will introduce relevant notation, and derive a variational formulation of System (1)–(3) which is suitable for a finite element approximation. Finally we present the fully discrete method and present the main result, i.e. preservation of total energy conservation and entropy production. Section 3 illustrates the theoretical scheme with a suitable convergence test. In Section 4 we conclude the work and present an outlook for future work.
@@ 2 Notation And Main Result
Before we present out discretization method and main results in detail, let us briefly introduce our notation and main assumptions, and recall some basic facts.
@@
We require the following assumptions for this work.
  • (A1) The interface parameters $\gamma$ is a positive constant.
  • (A2) The viscosity function $\eta\equiv\eta(\phi,\theta)$ is strictly positive.
  • (A3) The diffusion matrix $\mathbf{L}\equiv\mathbf{L}(\rho,\nabla\rho,\theta)\in\mathbb{R}^{2d\times 2d}$ given by $\begin{pmatrix}\mathbf{L}_{11}&-\mathbf{L}_{12}\\ -\mathbf{L}_{12}&\mathbf{L}_{22}\end{pmatrix}$ is symmetric and strictly positive definite.
  • (A4) For the driving potential $\Psi(\cdot,\cdot):\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ we assume that for every fixed $\phi$ the potential $\Psi(\phi,\cdot):\mathbb{R}_{+}\to\mathbb{R}$ is concave and goes to infinity for $\theta\to 0$. Furthermore, for every fixed $\theta\in\mathbb{R}_{+}$ the potential $\Psi(\cdot,\theta)$ can be decomposed in a strictly convex and a strictly concave part, denoted by $\Psi_{vex},\Psi_{cav}.$
Variational Formulation: In this paragraph we will recast the system into a variational form which is directly usable for a conforming finite element discretisation. We will see that the correct formulation of the Korteweg stress $\boldsymbol{\sigma}$ will be crucial. In the isothermal case it is well-known that a reformulation of $\boldsymbol{\sigma}$ and the pressure $p$ in terms of $\phi,\mu,\pi$, where $\pi$ is a modified pressure, allows to deduce such variational formulations and we will provide an non-isothermal expansion which reveal the strong coupling of the temperature.
@@
A sufficiently regular solution $(\phi,e,\mathbf{u},p)$ of the system (1)–(3) fulfills the following variational formulation
$\langle\partial_{t}\phi,\psi\rangle$ $-\langle\phi\mathbf{u},\nabla\psi\rangle+\langle\mathbf{L}_{11}\nabla\mu-\mathbf{L}_{12}\nabla\theta,\nabla\psi\rangle=0,$
$\langle\mu,\xi\rangle$ $-\gamma\langle\nabla\phi,\nabla\xi\rangle-\langle\partial_{\phi}\Psi,\xi\rangle=0,$
$\langle\partial_{t}e,w\rangle$ $+\langle\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta,\nabla w\rangle-\langle\eta|\mathrm{D}\mathbf{u}|^{2},w\rangle$
@@
$\langle\partial_{t}\mathbf{u},\mathbf{v}\rangle$ $+\mathbf{c}_{skw}(\mathbf{u},\mathbf{u},\mathbf{v})+\langle\eta\mathrm{D}\mathbf{u},\mathrm{D}\mathbf{v}\rangle-\langle\pi,\operatorname{div}(\mathbf{v})\rangle$
@@
00 0 $=\langle\operatorname{div}(\mathbf{u}),q\rangle$
for sufficiently regular test functions $\psi,\xi,\mathbf{v},w,q$ and $\pi:=p+e-\frac{s+\phi\mu}{\theta}$.
@@
The main part of the procedure is standard, i.e. multiplication by test functions, integration by parts. For the Korteweg stress we expand as follows
@@
The Korteweg stress appears in the internal energy equation as well as the Navier-Stokes equation. In weak form we find
@@
This variational formulation allows to deduce the thermodynamics quantities immediately by inserting suitable test function.
Theorem 2.1
For a sufficiently regular solution $(\phi,\mu,\theta,\mathbf{u},\pi)$ of (1) – (3) conservation of mass and total energy as well as entropy production holds, i.e.
@@
$\langle\partial_{t}s(\phi,\theta),1\rangle$ $=\|\sqrt{\eta\theta}\mathrm{D}\mathbf{u}\|_{0}^{2}+\langle(\nabla\mu,\nabla\theta)^{\top},\mathbf{L}(\nabla\mu,\nabla\theta)^{\top}\rangle\geq 0.$
@@
Conservation of mass follows immediately by inserting $\psi=1$. Conservation of total energy is obtain by inserting $w=1$, $\mathbf{v}=\mathbf{u}$ $q=\pi$, using the skew-symmetry $\mathbf{c}_{skw}(\mathbf{u},\mathbf{u},\mathbf{u})=0$, i.e.
@@
$\quad+\langle\tfrac{\phi}{\theta}\nabla\mu+\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{u}\rangle+\langle s+\phi\mu,\mathbf{u}\cdot\nabla\tfrac{1}{\theta}\rangle$
$=0.$
@@
$\langle\partial_{t}s,1\rangle$ $=\langle\partial_{t}e,\theta\rangle-\langle\partial_{t}\phi,\partial_{\phi}\Psi\rangle-\gamma\langle\nabla\phi,\nabla\partial_{t}\phi\rangle=\langle\partial_{t}e,\theta\rangle-\langle\mu,\partial_{t}\phi\rangle$
@@
$=\langle\mathbf{L}_{11}\nabla\mu-\mathbf{L}_{12}\nabla\theta,\nabla\mu\rangle-\langle\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta,\nabla\theta\rangle+\langle\eta|\mathrm{D}\mathbf{u}|^{2},\theta\rangle+$
@@
Time Discretization: We partition the time interval $[0,T]$ into uniformly with step size $\tau>0$ and introduce $\mathcal{I}_{\tau}:=\{t^{0}=0,t^{1}=\tau,\ldots,t^{n_{T}}=T\}$, where $n_{T}=\tfrac{T}{\tau}$ is the absolute number of time steps. We denote by $\Pi^{1}_{c}(\mathcal{I}_{\tau};X),\Pi^{0}(\mathcal{I}_{\tau};X)$ the spaces of continuous piece-wise linear and piece-wise constant functions on $\mathcal{I}_{\tau}$ with values in the space or set $X$. By $g^{n+1},g^{n},g^{n+1/2}$ we denote the new, the old time level and the midpoint approximation of $g$, i.e. $(g^{n+1}+g^{n})/2$. We introduce the time difference and the discrete time derivative via
@@
Space Discretization: For the spacial discretisation we require that $\mathcal{T}_{h}$ is a geometrically conforming partition of $\Omega$ into simplices that can be extended periodically to periodic extensions of $\Omega$. We denote the space of continuous, piecewise linear and quadratic functions over $\mathcal{T}_{h}$ as well as the space of mean free and set of positive piecewise-linear functions over $\mathcal{T}_{h}$ via
@@
We introduce the abbreviation for the convex-concave splitting by
$\Psi(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1})=\Psi_{vex}(\phi_{h}^{n+1},\theta_{h}^{n+1})+\Psi_{cav}(\phi_{h}^{n},\theta_{h}^{n+1})$
@@
We then propose the fully discrete time-stepping method for the CHNST system.
@@
Let $(\phi_{h,0},\mathbf{u}_{h,0},\theta_{h,0})\in\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{V}_{h}^{+}$ be given. Find the functions $(\phi_{h},\mathbf{u}_{h},\theta_{h})\in\Pi^{1}_{c}(\mathcal{I}_{\tau};\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{V}_{h}^{+})$ and $(\mu_{h},\pi_{h})\in\Pi^{0}(\mathcal{I}_{\tau};\mathcal{V}_{h}\times\mathcal{Q}_{h})$ such that
@@
$\langle\mu_{h}^{n+1},\xi_{h}\rangle$ $-\gamma\langle\nabla\phi_{h}^{n+1},\nabla\xi_{h}\rangle-\langle\partial_{\phi}\Psi(\phi^{n+1}_{h},\phi^{n}_{h},\theta^{n+1}_{h}),\xi_{h}\rangle=0,$
@@
holds for $(\psi_{h},\xi_{h},\mathbf{v}_{h},q_{h},w_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{Q}_{h}\times\mathcal{V}_{h}^{+}$. Here $g^{*}$ denotes an evaluation of $g$ at any $t\in\{t^{n},t^{n+1}\}$, but all terms have to be evaluated at the same point in time.
@@
$\langle\phi^{n+1}-\phi^{0},1\rangle=0,\qquad\langle\tfrac{1}{2}|\mathbf{u}^{n+1}_{h}|^{2}+e(\phi_{h}^{n+1},\theta_{h}^{n+1})-\tfrac{1}{2}|\mathbf{u}^{0}_{h}|^{2}-e(\phi_{h}^{0},\theta_{h}^{0},1\rangle=0,$
@@
For total energy conservation we insert $\mathbf{v}=\mathbf{u}^{n+1/2},w=1$ together with cancellation and the skew-symmetric of $\mathbf{c}_{skew}(\mathbf{u}_{h}^{*},\mathbf{u}_{h}^{n+1/2},\mathbf{u}_{h}^{n+1/2})=0$ we obtain
@@
Adding ±$\pm\partial_{\phi}\Psi(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1})d^{n+1}\phi_{h}$ and insertion of $\xi_{h}=d^{n+1}\phi_{h}$ yields
@@
Insert $\psi_{h}=-\tau\mu_{h}^{n+1},w_{h}=\tau\theta_{h}^{n+1}$ into the discrete formulation yields by the same calculations as in the continuous case for the entropy production equation. Finally, for the numerical dissipation we add ±$\pm\Psi_{h}(\phi_{h}^{n},\theta_{h}^{n+1})$ which yields
@@
$+(\partial_{\phi\phi}\Psi_{vex}(\xi^{1}_{h},\theta^{n+1}_{h})-\partial_{\phi\phi}\Psi_{cav}(\xi^{2}_{h},\theta^{n+1}_{h}))(d^{n+1}\phi_{h}^{n+1})^{2},$
where $\xi^{1}_{h},\xi^{2}_{h}$ are convex combinations of $\phi_{h}^{n+1},\phi_{h}^{n}$ and $\xi^{3}_{h}$ is a convex combination of $\theta_{h}^{n+1},\theta_{h}^{n}$. Using the structural assumptions on the potential $\Psi$, cf. (A4), we see that $\mathcal{D}_{num}\geq 0$ follows directly.

### v3

@@ Abstract
In this work we propose and analyse a structure-preserving approximation of the non-isothermal Cahn-Hilliard-Navier-Stokes system using conforming finite elements in space and implicit time discretisation with convex-concave splitting. The system is first reformulated into a variational form that reveals the equations’ structure and is then used in the subsequent approximation.
@@ 1 Introduction
The non-isothermal Cahn-Hilliard-Navier-Stokes (CHNST) system has increasingly gained attention for the investigation of different phenomena ranging from two-phase flows to fluid-phase-coupled interactions both of which have relevant importance in scientific and industrial applications, such as additive manufacturing and inkjet printing van2017binary ; Yang2020 ; dadvand2021advected . For instance, the modelling and simulation of powder bed fusion additive manufacturing (PBF-AM) process employ the non-isothermal CHNST to depict coupled processes in PBF-AM such as fluid-phase interaction, melt flow dynamics and heat transfer Yang2020 . For the CHNST system, we consider the following system of partial differential equations
@@
The above system is complemented by periodic boundary conditions and initial conditions. In this context $\phi$ denotes a conserved phase-field variable, $\mathbf{u}$ is the flow velocity, $\theta$ is the inverse temperature and $e\equiv e(\phi,\theta)$ is the internal energy. To close the system we consider the Helmholtz free energy density and the Korteweg stress given by
@@
$\tilde{\Psi}(\phi,\theta):=\Psi(\phi,\theta)+\tfrac{\gamma}{2}|\nabla\phi|^{2},\qquad\boldsymbol{\sigma}:=\tfrac{\gamma}{\theta}\nabla\phi\otimes\nabla\phi.$
@@
From this, one can compute the internal energy and entropy according to Alt1990 ; Alt1992 , i.e. $e=\partial_{\theta}\tilde{\Psi}$ and $\hat{s}(e(\theta,\phi),\phi)=s(\theta,\phi)=\theta e-\tilde{\Psi}$. Note that $\Psi(\phi,\theta)$ denotes the gradient-free part of the Helmholtz free energy density. A generic choice can be found in Section 3.
@@
The above system is already formulated in the inverse temperature $\theta$ which is more convenient for a finite element discretisation. The transformation from the original temperature to the inverse temperature can be found by Pawlow and Alt Alt1992 , and brunk2023variational for more complicated non-isothermal Cahn-Hilliard-Allen-Cahn system. The general derivation of non-isothermal phase-field models from thermodynamic sound principles is more delicate and we refer to Charach1998 ; Fabrizio2006 ; Pawlow2016 for several different modelling approaches.
@@
For this type of systems in Guo_2015 ; Sun2020 the authors consider finite differences in space together with an Energy Quadratisation (EQ) ansatz. This renders the driving functional quadratic, which facilitates standard time discretisation methods. However, the price of such an ansatz is the relaxation of entropy, which in some cases results in a drift-off from the original entropy. The isothermal version of the above system, i.e. the Cahn-Hilliard-Navier-Stokes system, is called model H introduced by Hohenberg and Halperin Hohenberg . This model is employed in the context of two-phase flow, especially in phase separation. In this case, there are many well-established techniques to solve the system with finite elements Diegel2017 ; Feng06 ; Han2015 and finite differences Gong2018 ; ChenZhao2020 ; LiShen2022 in space.
@@
In this work, we will provide a corresponding fully discrete method for the Cahn-Hilliard-Navier-Stokes system using standard finite elements in space and suitable explicit and implicit time-stepping. The discretisation can be seen as a natural extension of the well-known convex-concave splitting for the isothermal Cahn-Hilliard-Navier-Stokes system to the non-isothermal case.
@@
The work is structured as follows. In Section 2 we will introduce relevant notation, and derive a variational formulation of system (1)–(3) which is suitable for a finite element approximation. Afterwards, we present the fully discrete method and present the main result, i.e. preservation of total energy conservation and entropy production. Section 3 illustrates the theoretical scheme with a suitable convergence test. In Section 4 we conclude the work and present an outlook for future research.
@@ 2 Notation And Main Result
Before we present our discretization method and main results in detail, let us briefly introduce our notation and main assumptions, and recall some basic facts.
@@
In the following proposition, we collect the necessary assumptions applied in the whole work.
Proposition 1
We require the following assumptions:
  1. (A1) The interface parameter $\gamma$ is a positive constant.
  2. (A2) The viscosity function $\eta\equiv\eta(\phi,\theta)$ is strictly positive.
  3. (A3) The diffusion matrix $\mathbf{L}\equiv\mathbf{L}(\rho,\nabla\rho,\theta)\in\mathbb{R}^{2d\times 2d}$ given by $\begin{pmatrix}\mathbf{L}_{11}&-\mathbf{L}_{12}\\ -\mathbf{L}_{12}&\mathbf{L}_{22}\end{pmatrix}$ is symmetric and strictly positive definite. Note that also $\mathbf{L}_{12}$ is assumed to be symmetric.
  4. (A4) For the driving potential $\Psi(\cdot,\cdot):\mathbb{R}\times\mathbb{R}_{+}\to\mathbb{R}$ we assume that for every fixed $\phi$ the potential $\Psi(\phi,\cdot):\mathbb{R}_{+}\to\mathbb{R}$ is concave and goes to infinity for $\theta\to 0$. Furthermore, for every fixed $\theta\in\mathbb{R}_{+}$ the potential $\Psi(\cdot,\theta)$ can be decomposed in a strictly convex and a strictly concave part, denoted by $\Psi_{vex},\Psi_{cav}.$
Variational Formulation: In this paragraph, we will recast the system into a variational form which is directly usable for a conforming finite element discretisation. We will see that the correct formulation of the Korteweg stress $\boldsymbol{\sigma}$ will be crucial. In the isothermal case it is well-known that a reformulation of $\boldsymbol{\sigma}$ and the pressure $p$ in terms of $\phi,\mu,\pi$, where $\pi$ is a modified pressure, allows us to deduce such variational formulations and we will provide a non-isothermal expansion which reveals the strong coupling of the temperature.
To derive a suitable variational formulation, we assume that a sufficiently regular solution of system (1)–(3) exists at least on a short time interval. Let us fix ideas by assuming a classical solution.
Definition 1
We call a solution $(\phi,\theta,u,p)$ of (1)–(3) classical solution, if (1)–(3) holds pointwise for every $(t,x)\in(0,T]\times\Omega$ and the following regularity assumption holds
$\phi$ $\in C^{1}([0,T],C^{4}(\bar{\Omega})),$ $\mu$ $\in C^{0}([0,T],C^{2}(\bar{\Omega})),$
$\theta$ $\in C^{1}([0,T],C^{2}(\bar{\Omega})),$ $\mathbf{u}$ $\in C^{1}([0,T],C^{2}(\bar{\Omega})),$
$p$ $\in C^{0}([0,T],C^{4}(\bar{\Omega})).$
We note that such solutions might not exist at all or only exist under suitable assumptions for small times. If the following variational formulation is suitable to prove the existence of generalized solutions is so far open and a topic of future research.
@@
Every classical solution $(\phi,\mu,\theta,\mathbf{u},p)$ of the system (1)–(3) fulfills the following variational formulation
$\langle\partial_{t}\phi,\psi\rangle$ $-\langle\phi\mathbf{u},\nabla\psi\rangle+\langle\mathbf{L}_{11}\nabla\mu-\mathbf{L}_{12}\nabla\theta,\nabla\psi\rangle=0,$ (4)
$\langle\mu,\xi\rangle$ $-\gamma\langle\nabla\phi,\nabla\xi\rangle-\langle\partial_{\phi}\Psi,\xi\rangle=0,$ (5)
$\langle\partial_{t}e,w\rangle$ $+\langle\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta,\nabla w\rangle-\langle\eta|\mathrm{D}\mathbf{u}|^{2},w\rangle$ (6)
@@
$\langle\partial_{t}\mathbf{u},\mathbf{v}\rangle$ $+\mathbf{c}_{skw}(\mathbf{u},\mathbf{u},\mathbf{v})+\langle\eta\mathrm{D}\mathbf{u},\mathrm{D}\mathbf{v}\rangle-\langle\pi,\operatorname{div}(\mathbf{v})\rangle$ (7)
@@
00 0 $=\langle\operatorname{div}(\mathbf{u}),q\rangle$ (8)
for $\psi,\xi,w,\mathbf{v},q\in C^{1}(\bar{\Omega})$ and $\pi:=p+e-\frac{s+\phi\mu}{\theta}$.
@@
The main part of the procedure is standard, i.e. multiplication by test functions, integration by parts. The Korteweg stress can be expanded as follows
@@
The Korteweg stress appears in the internal energy equation as well as the Navier-Stokes equation. To derive the weak form (4)-(8) we compute
@@
Note that due to the special structure of the Navier-Stokes equation, we can redefine the pressure i.e. switching from $p$ to $\pi$ as the main variable. For smooth solutions, $p$ and $\pi$ are both regular. This variational formulation allows us to deduce the thermodynamics quantities immediately by inserting suitable test functions.
Lemma 2
For a classical solution $(\phi,\mu,\theta,\mathbf{u},\pi)$ of (1) – (3), cf. Definition 1, the conservation of mass and total energy as well as entropy production holds, i.e.
@@
$\langle\partial_{t}s(\phi,\theta),1\rangle$ $=\|\sqrt{\eta\theta}\mathrm{D}\mathbf{u}\|_{0}^{2}+\langle(\nabla\mu,\nabla\theta)^{\top},\mathbf{L}(\nabla\mu,\nabla\theta)^{\top}\rangle=:\mathcal{D}_{\theta,\mathbf{L}}(\mathrm{D}\mathbf{u},\nabla\mu,\nabla\theta)\geq 0.$
@@
Conservation of mass follows immediately by inserting $\psi=1$. Conservation of total energy is obtain by inserting $w=1$, $\mathbf{v}=\mathbf{u},$ $q=\pi$, using the skew-symmetry $\mathbf{c}_{skw}(\mathbf{u},\mathbf{u},\mathbf{u})=0$, i.e.
@@
$\quad+\langle\tfrac{\phi}{\theta}\nabla\mu+\boldsymbol{\sigma}\tfrac{\nabla\theta}{\theta},\mathbf{u}\rangle+\langle s+\phi\mu,\mathbf{u}\cdot\nabla\tfrac{1}{\theta}\rangle=0.$
@@
$\langle\partial_{t}s,1\rangle$ $=\langle\partial_{t}e,\theta\rangle-\langle\partial_{t}\phi,\partial_{\phi}\Psi\rangle-\gamma\langle\nabla\phi,\nabla\partial_{t}\phi\rangle=\langle\partial_{t}e,\theta\rangle-\langle\mu,\partial_{t}\phi\rangle$ (9)
@@
$=\langle\mathbf{L}_{11}\nabla\mu-\mathbf{L}_{12}\nabla\theta,\nabla\mu\rangle-\langle\mathbf{L}_{12}\nabla\mu-\mathbf{L}_{22}\nabla\theta,\nabla\theta\rangle+\langle\eta|\mathrm{D}\mathbf{u}|^{2},\theta\rangle$
@@
Time Discretization: We partition the time interval $[0,T]$ into uniform sub-intervals with step size $\tau>0$ and introduce $\mathcal{I}_{\tau}:=\{t^{0}=0,t^{1}=\tau,\ldots,t^{n_{T}}=T\}$, where $n_{T}=\tfrac{T}{\tau}$ is the absolute number of time steps. We denote by $\Pi^{1}_{c}(\mathcal{I}_{\tau};X),\Pi^{0}(\mathcal{I}_{\tau};X)$ the spaces of continuous piecewise linear and piecewise constant functions on $\mathcal{I}_{\tau}$ with values in the space or set $X$. By $g^{n+1},g^{n},g^{n+1/2}$ we denote the new, the old time level and the midpoint approximation of $g$, i.e. $(g^{n+1}+g^{n})/2$. We introduce the time difference and the discrete time derivative via
@@
Space Discretization: For the spatial discretisation we require that $\mathcal{T}_{h}$ is a geometrically conforming partition of $\Omega$ into simplices that can be extended periodically to periodic extensions of $\Omega$. We denote the space of continuous, piecewise linear and quadratic functions over $\mathcal{T}_{h}$ as well as the space of mean free and set of positive piecewise linear functions over $\mathcal{T}_{h}$ via
@@
We introduce the abbreviation for the convex-concave splitting, recall assumption (A4), cf. Proposition 1 by
$\Psi_{sp}(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1}):=\Psi_{vex}(\phi_{h}^{n+1},\theta_{h}^{n+1})+\Psi_{cav}(\phi_{h}^{n},\theta_{h}^{n+1})$
@@
We then propose a fully discrete time-stepping method for the CHNST system.
@@
Let $(\phi_{h}^{0},\mathbf{u}_{h}^{0},\theta_{h}^{0})\in\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{V}_{h}^{+}$ be given. Find the functions $(\phi_{h},\mathbf{u}_{h},\theta_{h})\in\Pi^{1}_{c}(\mathcal{I}_{\tau};\mathcal{V}_{h}\times\mathcal{X}_{h}^{d}\times\mathcal{V}_{h}^{+})$ and $(\mu_{h},\pi_{h})\in\Pi^{0}(\mathcal{I}_{\tau};\mathcal{V}_{h}\times\mathcal{Q}_{h})$ such that
@@
$\langle\mu_{h}^{n+1},\xi_{h}\rangle$ $-\gamma\langle\nabla\phi_{h}^{n+1},\nabla\xi_{h}\rangle-\langle\partial_{\phi}\Psi_{sp}(\phi^{n+1}_{h},\phi^{n}_{h},\theta^{n+1}_{h}),\xi_{h}\rangle=0,$
@@
holds for $(\psi_{h},\xi_{h},w_{h},\mathbf{v}_{h},q_{h})\in\mathcal{V}_{h}\times\mathcal{V}_{h}\times\mathcal{V}_{h}^{+}\times\mathcal{X}_{h}^{d}\times\mathcal{Q}_{h}$. Here $g^{*}$ denotes an evaluation of $g$ at any $t\in\{t^{n},t^{n+1}\}$, but all terms have to be evaluated at the same point in time.
@@
$\langle\phi^{n+1}_{h}-\phi^{0}_{h},1\rangle=0,\qquad\langle\tfrac{1}{2}|\mathbf{u}^{n+1}_{h}|^{2}+e(\phi_{h}^{n+1},\theta_{h}^{n+1})-\tfrac{1}{2}|\mathbf{u}^{0}_{h}|^{2}-e(\phi_{h}^{0},\theta_{h}^{0}),1\rangle=0,$
@@
Here $\xi^{1}_{h},\xi^{2}_{h}$ are convex combinations of $\phi_{h}^{n+1},\phi_{h}^{n}$ and $\xi^{3}_{h}$ is a convex combination of $\theta_{h}^{n+1},\theta_{h}^{n}$.
@@
For total energy conservation we insert $\mathbf{v}_{h}=\mathbf{u}_{h}^{n+1/2},w_{h}=1$ together with cancellation and the skew-symmetric of $\mathbf{c}_{skew}(\mathbf{u}_{h}^{*},\mathbf{u}_{h}^{n+1/2},\mathbf{u}_{h}^{n+1/2})=0$ we obtain
@@
Adding ±$\pm\partial_{\phi}\Psi_{sp}(\phi_{h}^{n+1},\phi_{h}^{n},\theta_{h}^{n+1})d^{n+1}\phi_{h}$ and insertion of $\xi_{h}=d^{n+1}\phi_{h}$ yield
@@
Insert $\psi_{h}=-\tau\mu_{h}^{n+1},w_{h}=\tau\theta_{h}^{n+1}$ into the discrete formulation we can mimic the same calculations as in the continuous case for the entropy production equation, cf. proof of Lemma 2, cf. (9). Hence, we obtain that
$\langle d^{n+1}e_{h},\theta_{h}^{n+1}\rangle-\langle\mu_{h}^{n+1},d^{n+1}\phi_{h}\rangle=\tau\mathcal{D}_{\theta^{n+1}_{h},\mathbf{L}^{*}}(\mathrm{D}\mathbf{u}^{n+1}_{h},\mu_{n}^{n+1},\theta_{h}^{n+1})$
Finally, for the numerical dissipation we add ±$\pm\Psi_{h}(\phi_{h}^{n},\theta_{h}^{n+1})$ which yields
@@
$+(\partial_{\phi\phi}\Psi_{vex}(\xi^{1}_{h},\theta^{n+1}_{h})-\partial_{\phi\phi}\Psi_{cav}(\xi^{2}_{h},\theta^{n+1}_{h}))(d^{n+1}\phi_{h}^{n+1})^{2}.$
Using the structural assumptions on the potential $\Psi$, cf. (A4) in Proposition 1, we see that $\mathcal{D}_{num}\geq 0$ follows directly.